# NAVINACT: Combining Navigation and Imitation Learning for Bootstrapping Reinforcement Learning

### v1

@@ 4.2 Simulation Experiments
Our evaluation suite includes three tasks from MetaWorld [37], each using sparse 0/1 task completion rewards at the end of each episode. These tasks are a subset of those evaluated in IBRL [28] and cover medium and hard difficulty levels as categorized in [35].
@@
Implementation of NAVINACT and Baselines:
NAVINACT employs Behavior Cloning (BC) for imitation learning (IL) and Twin Delayed Deep Deterministic Policy Gradient (TD3) for reinforcement learning (RL). A ResNet-18 vision encoder is used for BC, and random-shift image augmentation is applied to enhance RL performance. NAVINACT is benchmarked against three baselines: IBRL [28], RL-MN (RL with ModeNet and NavNet), and standard RL.
IBRL uses a pre-trained IL policy, $\mu_{\psi}$, to aid exploration during RL and to assist in target value estimation during Temporal Difference (TD) learning. Likewise, NAVINACT also uses IL policy within the InteractNet, that speeds up RL during object interaction. RL with ModeNet and NavNet, which we refer as RL-MN, incorporates ModeNet and NavNet to accelerate exploration by predicting waypoints, reducing the number of exploration steps needed in RL. Standard RL utilizes TD3 but does not use any IL or hierarchical mode prediction components.
@@ 4.3 Results
Following the training of NAVINACT and the baselines, we evaluate its performance on unseen environments and compare it against baseline methods. Performance is assessed using training and evaluation scores, which indicate sample efficiency and generalizability. Fig. 8 illustrates the comparative results for the all the metaworld environment, including training and evaluation scores for various initial positions.
@@
(d) Training with random inital Objects position.
(e) Training with random inital Objects position.
(f) Training with random inital Objects position.
@@
Figure 8: Performance Evaluation of NAVINACT Across MetaWorld Tasks. This figure presents the training and evaluation results for three challenging MetaWorld environments: Assembly, BoxClose, and CoffeePush. The top row depicts the environments in their initial configurations. The middle row shows the training performance of NAVINACT under conditions where object positions are randomized. The bottom two rows illustrate the evaluation performance under different randomized scenarios: varying initial object positions and combined randomization of both object and gripper positions. The results demonstrate NAVINACT’s robustness and adaptability to varying task complexities, maintaining high performance across different training and evaluation conditions.
Both NAVINACT and IBRL utilize a trained BC policy, with NAVINACT achieving a higher evaluation score of 0.38 in the assembly environment, while RL-MN and RL do not incorporate this policy. Our evaluation is focused on 3 key metrics — training score for the simulation tasks with randomization with respect to objects position, evaluation score with randomization with respect to objects position and evaluation score with randomization with respect to random initial gripper (as shown in Fig. 6 as well as object positions. The performance of NAVINACT, IBRL, RL-MN, and standard RL is compared across the Meta-World environments over 60k timesteps and 7 seeds.
NAVINACT outperforms baselines across all tasks in terms of both sample efficiency and final performance, solving all tasks within 30K samples.
4.3.1 Does NAVINACT Lead to Better Sample Efficiency?
To assess sample efficiency, we compare the training curves of NAVINACT with those of IBRL, RL-MN, and RL across various MetaWorld tasks. These tasks represent different levels of complexity, categorized into medium, hard, and very hard tiers. NAVINACT demonstrates better sample efficiency, as evidenced by faster convergence and higher success rates with fewer interaction steps as shown in Fig. 8(d),8(e) and 8(f). This improvement is attributed to NAVINACT’s mode prediction, which effectively reduces task complexity by narrowing down the action space at each step. NAVINACT consistently achieves an impressive success rate of 85±5% during training and learns quickly within 30k sample steps.
@@ 4.3.1 Does NAVINACT Lead to Better Sample Efficiency?
4.3.2 Does NAVINACT Generalize to Unseen Conditions?
To evaluate generalization, we test NAVINACT in conditions not included in the training data, such as variations in object positions, and initial gripper position. NAVINACT shows strong generalization capabilities, significantly outperforming IBRL and pure RL methods in unseen scenarios as indicated in Fig. 8(g),8(h),8(i),8(j),8(k) and 8(l). The mode prediction network plays a crucial role in adapting to new conditions by dynamically adjusting the operational mode, which is not achievable by methods solely relying on IL. Our algorithm maintains a nearly equivalent rate of 80±5% in evaluation phases across unseen environments. NAVINACT consistently outperforms other baselines in MetaWorld tasks, demonstrating its robustness to variability and changes in the environment.
We see that, NAVINACT matches or surpasses other baselines in Meta-World. The results indicate that, NAVINACT outperforms the other three methods in both training and evaluation. Notably, as randomization increases, the performance of IBRL, RL-MN, and standard RL decreases, while NAVINACT maintains consistent performance and demonstrates the robustness to variability and changes in environment.
@@ 4.4 Real-world Experimental Setup
Lift: The objective is to pick up a foam block. The initial location of the block is randomized over roughly $22cm\times(20cm-25cm)$ trapezoid, which covers the entire area visible from the wrist-camera when the robot is at the home position. See Fig.10(a), Fig. 10(b) for reference. We collect 10 demonstrations for this task due to its simplicity. It uses wrist-camera images as observations. We detect whether the gripper is holding the block by checking if the gripper width is static and the desired gripper width is smaller than the actual gripper width. The success detector returns 1 if the end effector has move upward by at least 2cm while holding the block. The maximum episode length is 120.
Pick-and-Place: The objective is to pick a soft toy on a wooden block. The initial location of the soft toy and the block is in a fixed position. The soft toy is initialized so that its visible from the wrsit camera view, see Fig.10(c). We use 20 prior demonstrations. This task uses environment camera images as observations as well because the wrist-camera loses sight of the block after picking up the soft-toy. We detect whether the gripper is holding the oft-toy by checking if the gripper width is static and the desired gripper width is smaller than the actual gripper width. The success returns 1 when the gripper width is static, the end-effector position are near the threshold of desired block position. The maximum episode legnth is 300.
@@
Figure 10: Top: Illustrations of lift task and the variation in the initialization of the task, left (fixed) right (randomized block position). Bottom: Performance evaluation on the lift task under both fixed and randomized conditions. The training curves depict the success rate over time, demonstrating NAVINACT’s effective learning and adaptability in varying scenarios.
As the primary goal of our real-world evaluations is to compare sample-efficiency and performance of various algorithms, we design rule-based success detectors and perform manual reset between episodes to ensure accurate reward and initial conditions. Note that sparse 0/1 reward from the success detector is the only source of reward.
@@ 4.5 Real-world Experiment Results
Does NAVINACT Work Effectively in Real-World Scenarios? Fig. 10 presents the training curves of NAVINACT compared to IBRL and across all different tasks. Each task is allowed a different interaction budget based on its difficulty, and the success rate is measured over each 1000-step interval during training. NAVINACT consistently learns faster than IBRL, and in the most challenging task, Pick-and-Place, it is the only method that surpasses behavior cloning (BC) under a 5K interaction budget.
In real-world evaluations, we assessed the final performance of each method using identical initial conditions. In the Lift task, NAVINACT achieves the highest success rate in both regular and “Hard Eval” settings, where the block’s initial position is more challenging. Even when the task difficulty increases, NAVINACT maintains a high success of 90% rate, demonstrating its robustness to distributional shifts and unseen scenarios, unlike BC, which fails in harder settings due to its reliance on the static training dataset as shown in Fig. 10.
How does NAVINACT perform for long-horizon tasks? The Pick-and-Place task, the most difficult due to two stage task manipulation, further highlights NAVINACT’s advantages as displayed in Fig. 10. While BC also doesn’t performs well due to suboptimal demostration, other method IBRL fail to match its success rate. While BC achieves a limited success rate of 20%, NAVINACT surpasses this by reaching a 75% success rate, even in complex scenarios where small errors can lead to significant failures.
@@ 5.1 Conclusion
In this paper, we presented NAVINACT, a hierarchical policy framework that combines NavNet for waypoint planning, ModeNet for dynamic mode switching, and InteractNet for fine-grained manipulation. By integrating imitation learning (IL) with reinforcement learning (RL), NAVINACT effectively addresses challenges of sample efficiency and distributional shift, outperforming existing methods. Through extensive simulations and real-world experiments, we demonstrated NAVINACT’s superior performance, generalizability, and adaptability across various tasks. Our approach learns faster and requires fewer interaction steps due to the innovative integration of mode and waypoint prediction, making it a scalable and robust solution for complex robotic manipulation in real-world scenarios.
@@ 5.2 Limitation and Future Work
While NAVINACT demonstrates significant advancements in sample efficiency and generalizability, one limitation of our approach is the reliance on collecting high-quality demonstration data, which can be time-consuming and expensive. This dependence on expert-generated data constrains the scalability and applicability of our framework across a wider range of tasks and environments. Future work could focus on developing methods that allow non-experts to collect usable data quickly and efficiently, potentially through resources such as videos in the wild.

### v2

@@ 4.2 Simulation Experiments
Our evaluation suite includes three tasks from MetaWorld [37], each using sparse 0/1 task completion rewards at the end of each episode. These tasks cover medium and hard difficulty levels as categorized in [35].
@@
Implementation of PLANRL and Baselines:
PLANRL employs Behavior Cloning (BC) for imitation learning (IL) and Twin Delayed Deep Deterministic Policy Gradient (TD3) for reinforcement learning (RL). A ResNet-18 vision encoder is used for BC, and random-shift image augmentation is applied to enhance RL performance. PLANRL is benchmarked against three baselines: IBRL [28], RL-MN (RL with ModeNet and NavNet), and standard RL.
IBRL uses a pre-trained IL policy, $\mu_{\psi}$, to aid exploration during RL and to assist in target value estimation during Temporal Difference (TD) learning. Likewise, PLANRL also uses IL policy within the InteractNet, that speeds up RL during object interaction. RL with ModeNet and NavNet, which we refer as RL-MN, incorporates ModeNet and NavNet to accelerate exploration by predicting waypoints, reducing the number of exploration steps needed in RL. Standard RL utilizes TD3 but does not use any IL or hierarchical mode prediction components.
@@ 4.3 Results
Following the training of PLANRL and the baselines, we evaluate its performance on unseen configuration of object positions in the environments and compare it against baseline methods. Performance is assessed using training and evaluation scores, which indicate sample efficiency and generalizability. Fig. 8 illustrates the comparative results for all the metaworld environments, including training and evaluation scores for various initial positions.
@@
(d) Training with random initial Objects position.
(e) Training with random initial Objects position.
(f) Training with random initial Objects position.
@@
Figure 8: Performance Evaluation of PLANRL Across MetaWorld Tasks. This figure presents the training and evaluation results for three challenging MetaWorld environments: (a) Assembly, (b) BoxClose, and (c) CoffeePush. The top row depicts these environments in their initial configurations. The middle row (d), (e) and (f) shows the training performance of PLANRL under conditions where object positions are randomized within environments (a), (b), and (c). The bottom two rows illustrate the evaluation performance under different randomized scenarios: varying initial object positions and combined randomization of both object and gripper positions for each environment. The results demonstrate PLANRL’s robustness and adaptability to varying task complexities, maintaining high performance across different training and evaluation conditions.
Both PLANRL and IBRL utilize a trained BC policy, with PLANRL achieving a higher evaluation score of 0.38 in the assembly environment, while RL-MN and RL do not incorporate this policy. Our evaluation is focused on 3 key metrics — training score for the simulation tasks with randomization with respect to object position, evaluation score with randomization with respect to object position, and evaluation score with randomization with respect to random initial gripper (as shown in Fig. 6) as well as object positions. The performance of PLANRL, IBRL, RL-MN, and standard RL is compared across the Meta-World environments over 60k timesteps and 7 seeds.
PLANRL outperforms baselines across all tasks in terms of both sample efficiency and final performance, solving all tasks within 30K samples.
4.3.1 Does PLANRL Lead to Better Sample Efficiency?
To assess sample efficiency, we compare the training curves of PLANRL with those of IBRL, RL-MN, and RL across various MetaWorld tasks. These tasks represent different levels of complexity, categorized into medium, hard, and very hard tiers. PLANRL demonstrates better sample efficiency, as evidenced by faster convergence and higher success rates with fewer interaction steps as shown in Fig. 8(d),8(e) and 8(f). This improvement is attributed to PLANRL’s mode prediction, which effectively reduces task complexity by narrowing down the action space at each step. PLANRL consistently achieves an impressive success rate of 85±5% during training and learns quickly within 30k sample steps.
@@ 4.3.1 Does PLANRL Lead to Better Sample Efficiency?
4.3.2 Does PLANRL Generalize to Unseen Conditions?
To evaluate generalization, we test PLANRL in conditions not included in the training data, such as variations in object positions, and initial gripper position. PLANRL shows strong generalization capabilities, significantly outperforming IBRL and pure RL methods in unseen scenarios as indicated in Fig. 8(g),8(h),8(i),8(j),8(k) and 8(l). The mode prediction network plays a crucial role in adapting to new conditions by dynamically adjusting the operational mode, which is not achievable by methods solely relying on IL. Our algorithm maintains a nearly equivalent rate of 80±5% in evaluation phases across unseen configurations of object position in the environments. We see that, PLANRL matches or surpasses other baselines in Meta-World. Notably, as randomization increases, the performance of IBRL, RL-MN, and standard RL decreases, while PLANRL maintains consistent performance and demonstrates robustness to variability and changes in the environment.
@@ 4.4 Real-world Experimental Setup
Lift: The objective is to pick up a foam block. The initial location of the block is randomized over roughly $22cm\times(20cm-25cm)$ trapezoid, which covers the entire area visible from the wrist-camera when the robot is at the home position. See Fig.10(a), Fig. 10(b) for reference. We collect 10 demonstrations for this task due to its simplicity. It uses wrist-camera images as observations. We detect whether the gripper is holding the block by checking if the gripper width is static and the desired gripper width is smaller than the actual gripper width. The success detector returns 1 if the end effector has moved upward by at least 2cm while holding the block. The maximum episode length is 120.
Pick-and-Place: The objective is to pick a soft toy on a wooden block. The initial location of the soft toy and the block is in a fixed position. The soft toy is initialized so that it is visible from the wrist camera view, see Fig.10(c). We use 20 prior demonstrations. This task uses environment camera images as observations as well because the wrist-camera loses sight of the block after picking up the soft toy. We detect whether the gripper is holding the soft toy by checking if the gripper width is static and the desired gripper width is smaller than the actual gripper width. The success returns 1 when the gripper width is static, the end-effector position is near the threshold of desired block position. The maximum episode length is 300.
@@
Figure 10: Top: Illustrations of lift task and the variation in the initialization of the task, left (fixed) right (randomized block position). Bottom: Performance evaluation on the lift task under both fixed and randomized conditions. The training curves depict the success rate over time, demonstrating PLANRL’s effective learning and adaptability in varying scenarios.
As the primary goal of our real-world evaluations is to compare sample efficiency and performance of various algorithms, we design rule-based success detectors and perform manual reset between episodes to ensure accurate reward and initial conditions. Note that the sparse 0/1 reward from the success detector is the only source of reward.
@@ 4.5 Real-world Experiment Results
Does PLANRL Work Effectively in Real-World Scenarios? Fig. 10 presents the training curves of PLANRL compared to IBRL and across all different tasks. Each task is allowed different interactions based on its difficulty, and the success rate is measured over each 1000-step interval during training. PLANRL consistently learns faster than IBRL, and in the challenging task, Pick-and-Place, it is the only method that surpasses behavior cloning (BC) under 5K interactions.
In real-world evaluations, we assessed the final performance of each method using identical initial conditions. In the Lift task, PLANRL achieves the highest success rate in both regular and “Hard Eval” settings, where the block’s initial position is more challenging. Even when the task difficulty increases, PLANRL maintains a high success of 90% rate, demonstrating its robustness to distributional shifts and unseen scenarios, unlike BC, which fails in harder settings due to its reliance on the static training dataset as shown in Fig. 10.
How does PLANRL perform for long-horizon tasks? The Pick-and-Place task, the most difficult due to two-stage task manipulation, further highlights PLANRL’s advantages as displayed in Fig. 10. While BC also doesn’t perform well due to suboptimal demonstration, other methods IBRL fail to match its success rate. While BC achieves a limited success rate of 20%, PLANRL surpasses this by reaching a 75% success rate, even in complex scenarios where small errors can lead to significant failures.
@@ 5.1 Conclusion
In this paper, we presented PLANRL, a hierarchical policy framework that combines NavNet for waypoint motion-planning, ModeNet for dynamic mode switching, and InteractNet for fine-grained manipulation. By integrating imitation learning (IL) with reinforcement learning (RL), PLANRL effectively addresses challenges of sample efficiency and distributional shift, outperforming existing methods. Through extensive simulations and real-world experiments, we demonstrated PLANRL’s superior performance, generalizability, and adaptability across various tasks. Our approach learns faster and requires fewer interaction steps due to the integration of mode and waypoint prediction, making it a scalable and robust solution for complex robotic manipulation in real-world scenarios.
@@ 5.2 Limitation and Future Work
While PLANRL demonstrates significant advancements in sample efficiency and generalizability, one limitation of our approach is the reliance on collecting high-quality demonstration data, which can be time-consuming and expensive. Future work could focus on developing methods that allow non-experts to collect usable data quickly and efficiently, potentially through resources such as videos in the wild or collecting demonstrations using sketches.